# Maximum Preference Optimization with Importance Sampling

### 31 Dec 2023 v2

@@ 4.2 Off-policy Preference Learning under KL-regularation
$\max_{\pi^{s}_{\theta}}R(\pi^{p}_{\theta})-\beta\underset{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(y|x)]-\gamma\underset{x\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(x)]$ (31)
@@
$\mathcal{L}_{\text{MPO}}=\underset{\begin{subarray}{c}<x,y_{w},y_{l},\mathfrak{a},I>\sim\bar{\mathcal{D}}^{p}\\ \end{subarray}}{\mathbb{E}}[-I\pi^{p}_{\theta}(\mathfrak{a}|x)]+\beta\underset{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(y|x)]+\gamma\underset{y\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log{\pi^{s}_{\theta}(y|x)}].$ (32)

### 04 Jan 2024 v3

@@ 4.2 Off-policy Preference Learning under KL-regularation
$\max_{\pi^{s}_{\theta}}R(\pi^{p}_{\theta})+\beta\underset{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(y|x)]+\gamma\underset{x\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(x)]$ (31)
@@
$\mathcal{L}_{\text{MPO}}=\underset{\begin{subarray}{c}<x,y_{w},y_{l},\mathfrak{a},I>\sim\bar{\mathcal{D}}^{p}\\ \end{subarray}}{\mathbb{E}}[-I\pi^{p}_{\theta}(\mathfrak{a}|x)]-\beta\underset{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(y|x)]-\gamma\underset{x\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log{\pi^{s}_{\theta}(x)}].$ (32)